# Entropic balance with feedback control: information equalities and tight inequalities

### v2

@@ Abstract
These second-law variants have significantly advanced our understanding of measurement and feedback in physical systems, but crucial challenges remain. First, the tightness of the bounds on extractable work is not fully established  [^1] In a specific Brownian information engine, the bound involving unavailable information has been shown to become an equality [15, 52].. Second, computing quantities like the transfer entropy is challenging, both analytically and numerically, due to correlations between measurements  [^2] Thus, the transfer entropy has only been computed for very simple systems. For error-free measurements, recent work suggests that it can be estimated with moderately long chains [10].. This Letter addresses the above limitations, both fundamental and practical, in the state of the art.
@@
The controller updates its state in the infinitesimal interval $\tau_{\text{m}}^{(n)}\equiv(t_{n}^{-},t_{n}^{+})$, $c_{n}$ with conditional probability $\Theta(c_{n}|x_{n})$ at $t_{n}^{+}$. Then, $P(x,c,t_{n}^{+})=\Theta(c|x)P(x,t_{n})$, and $P(x,t)$ is continuous at $t_{n}$, $P(x,t_{n}^{+})=P(x,t_{n}^{-})$. For error-free measurement, $x_{n}$ univocally determines $c_{n}$: $\mathcal{D}$ splits into non-overlapping subsets $\mathcal{X}_{c}$, $\cup_{c}\mathcal{X}_{c}=\mathcal{D}$, and the controller switches on $V(x,c_{n})$ if $x_{n}\in\mathcal{X}_{c_{n}}$. Then, $\Theta(c|x)=\bm{1}_{c}(x)$, where $\bm{1}_{c}(x)$ is the indicator function of the interval $\mathcal{X}_{c}$. The numerical section shows a specific example of $\Theta(c|x)$ for measurements with errors.
@@
Equation (8) is our first key result. The relevant regime for information engines is $\left\langle W\right\rangle=\left\langle W\right\rangle_{\text{m}}+\left\langle W\right\rangle_{\text{FP}}<0$. Two limit cases are worth considering: (i) open-loop control; measurement is not used for control, thus $c_{n}$ is independent of $x_{n}$ and $S_{c}=S_{c|x}$; $\Delta_{\text{m}}\,I=-I_{-}<0$, (ii) error-free closed-loop control; $c_{n}$ is enslaved to $x_{n}$ and $S_{c|x}=0$; $\Delta_{\text{m}}\,I=S_{c|x}^{-}>0$. Thus, feedback-control—maybe imperfect—is essential for extracting work.
@@
Inspired by the experimental setup in Ref. [11], we consider a model with a potential $U(x)$ that can be inverted by the controller, V(x,c)=cU(x)V(x,c)=c\,U(x) with $c=\pm 1$. Specifically, we have considered the piecewise potential $U(x)=V_{0}(|x|-L/2)/L$ in the interval [L,L][-L,L] with periodic boundary conditions (left panel of Fig. 2888This choice preserves the physical image of a sinusoidal potential [11] while simplifying analytical calculations.. For the error-free case, $\Theta(c|x)=H(-c\,U(x))$, with $H(\cdot)$ being the Heaviside step function. When the particle position is measured with precision $\Delta x$, $\Theta(c|x)$ is proportional to the length of the interval $(x-\Delta x/2,x+\Delta x/2)$ with the “correct” sign of $U(x)$ (right panel of Fig. 2).
@@
Acknowledgments– We acknowledge financial support from Grant PID2024-155268NB-I00 funded by MICIU/AEI/10.13039/501100011033/ FEDER, UE. We also acknowledge support from Grant ProyExcel_00796 funded by Junta de Andalucía’s PAIDI 2020 program, and the applied research and innovation Project PPIT2024-31833, cofunded by EU–Ministerio de Hacienda y Función Pública–Fondos Europeos–Junta de Andalucía–Consejería de Universidad, Investigación e Innovación. NRP acknowledges support from the FPU programme through Grant FPU2021/01764.

### v3

@@ Abstract
These second-law variants have significantly advanced our understanding of measurement and feedback in physical systems, but crucial challenges remain. First, the tightness of the bounds on extractable work is not fully established  [^1] In a specific Brownian information engine, the bound involving the unavailable information has been shown to become an equality [15, 52].. Second, computing quantities like the transfer entropy is challenging, both analytically and numerically, due to correlations between measurements  [^2] Thus, the transfer entropy has only been computed for very simple systems. For error-free measurements, recent work suggests that it can be estimated with moderately long chains [10].. This Letter addresses the above limitations, both fundamental and practical, in the state of the art.
@@
The controller updates its state, in the infinitesimal interval $\tau_{\text{m}}^{(n)}\equiv(t_{n}^{-},t_{n}^{+})$, to $c_{n}$ with conditional probability $\Theta(c_{n}|x_{n})$ at $t_{n}^{+}$. Therefore, $P(x,c,t_{n}^{+})=\Theta(c|x)P(x,t_{n})$, and $P(x,t)$ is continuous at $t_{n}$, $P(x,t_{n}^{+})=P(x,t_{n}^{-})$. For error-free measurement, $x_{n}$ univocally determines $c_{n}$: $\mathcal{D}$ splits into non-overlapping subsets $\mathcal{X}_{c}$, $\cup_{c}\mathcal{X}_{c}=\mathcal{D}$, and the controller switches on $V(x,c_{n})$ if $x_{n}\in\mathcal{X}_{c_{n}}$. Then, $\Theta(c|x)=\bm{1}_{c}(x)$, where $\bm{1}_{c}(x)$ is the indicator function of the interval $\mathcal{X}_{c}$. The numerical section shows a specific example of $\Theta(c|x)$ for measurements with errors.
@@
Equation (8) is our first key result. The relevant regime for information engines is $\left\langle W\right\rangle=\left\langle W\right\rangle_{\text{m}}+\left\langle W\right\rangle_{\text{FP}}<0$. Two limit cases are worth considering: (i) open-loop control; measurement is not used for control, thus $c_{n}$ is independent of $x_{n}$ and $S_{c}=S_{c|x}^{+}$; $\Delta_{\text{m}}\,I=-I_{-}<0$, (ii) error-free closed-loop control; $c_{n}$ is enslaved to $x_{n}$ and $S_{c|x}^{+}=0$; $\Delta_{\text{m}}\,I=S_{c|x}^{-}>0$. Thus, feedback-control—maybe imperfect—is essential for extracting work.
@@
Inspired by the experimental setup in Ref. [11], we consider a model with a potential $U(x)$ that can be inverted by the controller, V(x,c)=cU(x)V(x,c)=c\,U(x) with $c=\pm 1$. Specifically, we have considered the piecewise potential $U(x)=V_{0}(|x|-L/2)/L$ in the interval [L,L][-L,L] with periodic boundary conditions (left panel of Fig. 2888This choice preserves the physical image of a sinusoidal potential [11] while simplifying analytical calculations.. For the error-free case, $\Theta(c|x)=H(-c\,U(x))$, with $H(\cdot)$ being the Heaviside step function. When the particle position is measured with precision $\Delta x$, $\Theta(c|x)$ is proportional to the length of the interval $(x-\Delta x,x+\Delta x)$ with the “correct” sign of $U(x)$ (right panel of Fig. 2).
@@
Acknowledgments– We acknowledge financial support from Grant PID2024-155268NB-I00 funded by MICIU/AEI/10.13039/501100011033/ FEDER, UE. We also acknowledge support from Grant ProyExcel_00796 funded by Junta de Andalucía’s PAIDI 2020 program, and the applied research and innovation Project PPIT2024-31833, cofunded by EU–Ministerio de Hacienda y Función Pública–Fondos Europeos–Junta de Andalucía–Consejería de Universidad, Investigación e Innovación. NRP acknowledges support from the FPU programme through Grant FPU21/01764.